Let G be a group and H be a non-empty finite subset of .G If ab∈H∀ b,a ∈ H then show that H ≤ .G Will the result remain true if H is not finite? Give reasons for your answer.
1
Expert's answer
2020-03-13T10:36:07-0400

We known that , A nonempty subset H of a group G is a subgroup if

"ab^{-1}\in H" when ever "a" and "b" in "H" .

In view of the above theorem ,we need only prove that "a^{-1}\in H \space whenever \space a\in H" .

If "a=e," then "a^{-1}=a" and we are done.If "a\neq e" consider the sequence. "a,a^2,a^2,........" By closure ,all of these elements belongs to "H" .Since "H" is finite ,not all of these elements are distinct.

Say "a^i=a^j \space and \space i>j." Then "a^{i-j}=e" and since "a\neq e, i-j>1." Thus ,"aa^{i-j-1}=a^{i-j}=e" and ,therefore,"a^{i-j-1}=a^{-1}." But "i-j-1\ge1" implies "a^{i-j-1}\in H" and we are done.

If H is not finite then we have a counter example.

Let "G=Z" (set of integer) group under addition and "H=N" (set of natural number).

Then "a+b\in N \space \forall \space a,b\in N"

But N is not a group under addition.